OpenStudy (anonymous):

rationalize the denominator: sqrt -16 / (1+i) + (6+3i)

OpenStudy (helder_edwin):

is this \[ \Large \frac{\sqrt{-16}}{1+i}+(6+3i) \] right?

OpenStudy (anonymous):

yes but the (6+3i) is is with the 1+i. its all on the botom

OpenStudy (lgbasallote):

\[\frac{\sqrt{-16}}{(1+i) + (6+3i)}\] \[\frac{4i}{1+i + 6 + 3i}\] \[\frac{4i}{7 + 4i}\] \[\frac{4i}{7+4i} \times \frac{7-4i}{7-4i}\]

OpenStudy (lgbasallote):

got it?

OpenStudy (anonymous):

could you break it down further please?